OpenStudy (samigupta8):

if tanθ1,tanθ2,tanθ3,tanθ4 are roots of equation x^4-x^3sin2β +x^2cos2β-xcosβ-sinβ=0 prove that tan(θ1+θ2+θ3+θ4)=cotβ

OpenStudy (anonymous):

u know the formula of tan(@1, @2, @3.......@n)= ?

OpenStudy (samigupta8):

tan(θ1+θ2+.......tanθn)=s1-s3+s5+......sn/1-s2+s4-.......+sn

OpenStudy (anonymous):

calm down dear....um still onlyn!! and just put the values in s1, s2 , s3, s4 in the formula

OpenStudy (samigupta8):

bt dat's wht i don't know what r the values of s1,s2,s3 and s4 here

OpenStudy (anonymous):

coeff of x^3= -(sum of roots taken one at time) x^2= +( two at a time) x = - ( 3 at a time ) const= +(product(i.e. four at a time))

OpenStudy (samigupta8):

yeh kya sum of roots of biquadatic equation nikal rhe ho aur phir product of roots e/a

OpenStudy (anonymous):

yeh jo likha h woh ek factorizable polynomial ke liye standard h

OpenStudy (samigupta8):

accha theek h usse toh aa gaya mera ans abbb

OpenStudy (samigupta8):

aur aap yeh btao ki aap ko mere msgs receive ho rhe h ya nahin

OpenStudy (anonymous):

pehle wale 4 to hue the...

OpenStudy (samigupta8):

accha

OpenStudy (anonymous):

try karunga reply karne ki..

OpenStudy (samigupta8):

aap try toh abhi dusri ques kr do filhaal msg ka toh sau din laga dena aaram se